Circular No. 6867

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

1SAX J0544.1–710 = RX J0544.1–7100 F. Haberl, W. Pietsch, and M. D. Filipovic, Max-Planck-Institut für

Extraterrestrische Physik, Garching, communicate: “The position of 1SAX J0544.1–710 (IAUC 6861) is consistent with that of RX J0544.1–7100. This ROSAT source is included in a list of variable x-ray sources resulting from a systematic study of the ROSAT PSPC pointed data of the LMC region (Haberl and Pietsch 1998, to be submitted to A.Ap.). The best ROSAT position is α = 5^{h}44^{m}06^{s}.3, δ = −71^{o}00′50′′ (equinox 2000.0; estimated 90-percent-confidence error 8′′). The source shows variations by at least a factor of 11 on timescales of years between four ROSAT PSPC and HRI observations. Given the variability and hard ROSAT spectrum, it is suggested as a high-mass x-ray binary candidate, probably a Be/x-ray binary in the LMC. A likely optical counterpart in the x-ray error circle is MACS J0544-710#002 from the Magellanic Catalogue of Stars (Tucholke et al. 1996, A.Ap. Suppl. 119, 91).”

NEW X-RAY BURSTER J. M. Muller, BeppoSAX Science Data Center, Rome, and Space Re-

search Organization of the Netherlands (SRON), Utrecht; M. J. S. Smith, BeppoSAX Science Operations Center (SOC), Rome, and SRON; G. D’Andreta, SOC; A. Bazzano and P. Ubertini, Istituto di Astrofisica Spaziale, CNR, Rome; and J. in ’t Zand and J. Heise, SRON, report: “An x-ray burst has been detected with Wide Field Camera 1 on BeppoSAX on Apr. 2.2714 UT from α = 18^{h}06^{m}48^{s}, δ = −24^{o}35′.9 (equinox 2000.0; error radius 4′). The burst had a peak intensity of 0.6 Crab in the energy band 2–9 keV and a duration of 15 s. No persistent emission was detected above an upper limit of 8 mCrab from Apr. 1.08 to 2.37 UT. The only catalogued x-ray source within the error circle is the transient 2S 1803–245 (Jernigan et al. 1978, Nature 272, 701), which is not known to be a burster.”

GRB 980329

Corrigendum. On IAUC 6866, last line, for ftp://astro.chicago.edu/

pub/astro/GRB/.” read ftp://astro.uchicago.edu/pub/astro/GRB/.”

1998 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green